OpenStudy (anonymous):

solve sin(2x) - cos(4x) = 0

OpenStudy (anonymous):

sin(2x)+2sin^2(x2)-1=0 (2sin(x)-1)(sin(2x)+1)=0 either sin(2x)=1/2 x=pi/12 or sin(2x)=-1 x=3pi/4